OpenStudy (anonymous):

how does 2cos5xsin5x turn into/equal sin10x?

OpenStudy (zzr0ck3r):

\(\sin(2*x)=2\sin(x)cos(x)\) so \(\sin(2*(5x))=2\sin(5x)\cos(5x)\) i.e. \(\sin(10x)=2\sin(5x)\cos(5x)\)

OpenStudy (anonymous):

ah... the application of the double angle formula confused me. thank you

OpenStudy (anonymous):

i didn't know how to use the double angle formula

OpenStudy (zzr0ck3r):

think of it anytime you have sin cosin next to each other for identities anyway....